OpenStudy (mr._scampy):

How would I go about solving for B, if T = (2B / B-1)?

OpenStudy (tkhunny):

First, you have written \(T = \dfrac{2B}{B} - 1\) 2B/(B-1) = \(\dfrac{2B}{B-1}\) Guessing that it is the latter, have you considered multiplying by "B-1" after noting that B = 1 is absolutely unacceptable?

OpenStudy (mr._scampy):

I apologize. Yes, I meant 2B/(B-1). I have tried multiplying both sides by "B-1", and I arrive at TB-T = 2B. And then I'm unsure of how to isolate B further.

OpenStudy (tkhunny):

Subtract TB